OpenStudy (daniel.ohearn1):

Integrate (cos(x))^4

OpenStudy (daniel.ohearn1):

Is there a way without the reduction formula?

ganeshie8 (ganeshie8):

A quick search gives you hundreds of results with different alternative methods

ganeshie8 (ganeshie8):

Here is the method that I like : \[(\cos x)^4 = \frac{1}{2^4}(e^{ix}+e^{-ix})^4 \] Expand above using binomial theorem and integrate term by term